OpenStudy (18jonea):

Find the angles in the triangle when a = 13, b = 12, and c = 11. Round your answers to the nearest tenth.

OpenStudy (18jonea):

@perl

OpenStudy (18jonea):

@Directrix

OpenStudy (anonymous):



OpenStudy (18jonea):

wait a = 68.7?

OpenStudy (anonymous):

it is actually angle A

OpenStudy (18jonea):

thats what i thought

OpenStudy (18jonea):

then b =59.3

OpenStudy (anonymous):

type error \[\cos A=\frac{ b^2+c^2-a^2 }{ 2bc }\]

OpenStudy (anonymous):

use sine formula to find other angle

OpenStudy (18jonea):

what other angle c= 52

OpenStudy (anonymous):

\[\frac{ c }{ \sin C }=\frac{ a }{ \sin A }\]

OpenStudy (18jonea):

i got 52 @surjithayer

OpenStudy (18jonea):

@Directrix

OpenStudy (18jonea):

@Vocaloid

OpenStudy (anonymous):

i have not solved other angles so i can't tell.

OpenStudy (18jonea):

can you double check mine please

OpenStudy (18jonea):

@surjithayer

OpenStudy (anonymous):

correct.